The length of timber cuts are normally distributed with a mean of 95inches. In a random sample of 45 boards, what is the probability that the mean of the sample will be between 94.8 inches and 95.8 inches?
1
Expert's answer
2020-09-29T18:10:38-0400

The following information about the mean and standard deviation has been provided:

μ=95, σ=0.52, n=45

We need to compute "\Pr(94.8 \le \bar X \le 95.8)"The corresponding z-values needed to be computed are:

"Z_1= \frac{X_1-\mu}{\sigma/\sqrt{n}} = \frac{ 94.8-95}{ 0.52/\sqrt{ 45}}= -2.5801\\Z_2 = \frac{X_2-\mu}{\sigma/\sqrt{n}} = \frac{ 95.8-95}{ 0.52/\sqrt{ 45}}= 10.3203"

Therefore, the following is obtained:

"Pr(94.8\u2264 \nX\n\u02c9\n \u226495.8)" "=Pr(\u22122.5801\u2264Z\u226410.3203)"

"=Pr(Z\u226410.3203)\u2212Pr(Z\u2264\u22122.5801)=1\u22120.0049=0.9951"